find the maximum(or minimum) of each function, also determine the x-intercepts of each function (if any). express your final answers as decimal number (rounded to 4 decimals if necessary)
a) f(x)=-x^2-4x-2
b) f(x)=-x^2+1x+3
I really dont understand these question, we went over them in class but I want to see them solved a Different way.

Answer by stanbon(75887)

find the maximum(or minimum) of each function, also determine the x-intercepts of each function (if any). express your final answers as decimal number (rounded to 4 decimals if necessary)
a) f(x)=-x^2-4x-2
You have a quadratic with a=-1,b=-4
The max or min is at x=-b/2a=-(-4/-2)=-2
Find f(-2)=-(-2)^2-4(-2)-2=-4+8=4
The max is at (-2,4)
-------------------
b) f(x)=-x^2+1x+3
a=-1,b=1
max or min at x=-1/(-2)=1/2
f(1/2)=(1/4)+(1/2)+3= 15/4
max at (1/2,15/4)
Cheers,
Stan H.


-----------------